OpenStudy (anonymous):

Chuck traveled 82 miles in the same time that Dana traveled 74 miles. If Chuck's rate of travel is 4mph more than Dana, and they traveled the same lenght of time at what speed does Chuck travel?

OpenStudy (anonymous):

let the speed of Chuck be x Then the speed of Dana is x-4 ( as Chuck is 4mph more than Dana) time=speed/ distance travelled x/82=(x-4)/74 By solving this equation, we can get x which is the speed of Chuck

OpenStudy (anonymous):

\[speed = \frac{distance}{time}\]\[time = \frac{speed}{distance}\]they traveled with the same length of time, chuck's speed is 4 more than dana, therefore dana's speed is 4 less than chuck, let the speed of chuck is x, then dana's speed is x - 4 \[t_{1} = t_{2}\]\[\frac{x}{82} = \frac{x - 4}{72}\] solve for x and we will get chuck's speed

OpenStudy (anonymous):

is it x=41?

OpenStudy (anonymous):

yes, x = 41

OpenStudy (anonymous):

x/82=(x-4)/74 74x=82(x-4) 74x=82x-328 8x=328 x=41 Therefore, yes!

OpenStudy (anonymous):

thanks guys ;)